• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name pking

    Log entry time 01:07:18 on October19,2009

    Entry number 295592

    keyword=Compton tripped; backgroud rates higer

    The Compton HV tripped at 01:02.

    The rates had clibed to 800-900 Hz/uA from ~700 with laser on,
    just before the trip. As seen in figure 1.

    Restored HV.

    FIGURE 1